OpenStudy (anonymous):

5x + 3(-2x + 1) = 8 Distribute, simplify, and simplify some more

OpenStudy (aroub):

5x-6x+3=8 5x-6x=-3+8 -x=5 x=-5

OpenStudy (anonymous):

5x-6x+3=8 5x-6x=-3+8 -x=5 x=-5